OpenStudy (anonymous):

Please help!! (4√5+7√15)/(6√15-√5) Rationalize the denominator.

OpenStudy (anonymous):

Multiply the numerator and denominator by the conjugate of the denominator, which would be (6√15+√5)

OpenStudy (anonymous):

(4√5+7√15)/(6√15-√5)*(6√15+√5)(6√15+√5)?

OpenStudy (anonymous):

yes basically. so top and bottom would be multiplied by the same thing

OpenStudy (anonymous):

(4√5+7√15)/(6√15-√5) * (6√15+√5)/(6√15+√5

OpenStudy (ranga):

\[\frac{ 4\sqrt{5} + 7\sqrt{15} }{ 6\sqrt{15} - \sqrt{5} } \times \frac{ 6\sqrt{15} + \sqrt{5} }{ 6\sqrt{15} + \sqrt{5} } \]

OpenStudy (anonymous):

ranga's visual is very helpful

OpenStudy (anonymous):

How do you multiply if √5 and √15 are not the same? I'm not understanding that part.

OpenStudy (anonymous):

Would I need to make them the same or leave them alone?

OpenStudy (ranga):

\[\sqrt{5} \times \sqrt{15} = \sqrt{5 \times 15} = \sqrt{75} = \sqrt{25 \times 3} = 5\sqrt{3}\]

OpenStudy (anonymous):

multiply what is under the radical

OpenStudy (ranga):

For rationalizing the denominator use the formula (a + b) x (a - b) = a^2 - b^2

OpenStudy (anonymous):

For 4√5 and 6√15 do I multiply the 4 and 6 together?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

24√75 would simplify to 120√3

OpenStudy (anonymous):

So for the numerator would be 24√75 +20 +42√225 +7√75 without being simplified?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

And the denominator would be (6√15)^2 (√5)^2 ?

OpenStudy (ranga):

with a minus sign in between

OpenStudy (anonymous):

(6√15)^2-(√5)^2

OpenStudy (ranga):

yes

OpenStudy (anonymous):

Then do I combine like terms in the numerator?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

31√75+20+42√225?

OpenStudy (ranga):

simplify the denominator too. You will get a nice round number without all the radicals. That is the whole idea behind multiplying by the conjugate pair.

OpenStudy (anonymous):

How do I do that exactly?

OpenStudy (anonymous):

well your denominator should have the 6√15^-√5^2. so with those numbers, you can make them simpler.

OpenStudy (ranga):

(6√15)^2-(√5)^2 = (36)(15) - (5)

OpenStudy (anonymous):

I'm not following. Sorry.

OpenStudy (ranga):

(ab)^2 = (a^2)(b^2) The radical sign here is the square root. If you take the square root of a number and then square it the square root and the square will cancel out and you will be left with the number. So if you square (square root of 15) you will be left with 15.

OpenStudy (anonymous):

Okay. I got it but I'm still not understanding the simplifying the numerator.

OpenStudy (ranga):

\[\sqrt{75} = \sqrt{25 \times 3} = 5\sqrt{3}\]

OpenStudy (anonymous):

So for 24√75 I would multiply the 5 and the 24 together?

OpenStudy (ranga):

yes

OpenStudy (anonymous):

So √225 would be 75√3?

OpenStudy (anonymous):

or just 15

OpenStudy (anonymous):

How did you get the 15?

OpenStudy (anonymous):

42√225 would be 42(15)

OpenStudy (anonymous):

the square root of 225 is 15

OpenStudy (anonymous):

120√3+20+15?

OpenStudy (anonymous):

120√3+20+(42)(15) + 35√3 should be in your numerator.

OpenStudy (ranga):

I have not double checked my numbers but I am getting \[\frac{ 634 + 155\sqrt{3} }{ 535 }\] If OCity gets the same that would be like double checking!

OpenStudy (anonymous):

I've got the same!

OpenStudy (ranga):

Cool!

OpenStudy (anonymous):

I get it. Thanks guys!

OpenStudy (anonymous):

you're welcome :)

OpenStudy (ranga):

yw

OpenStudy (anonymous):

There isn't a way to give you both a best response is there?

OpenStudy (anonymous):

no, go ahead and give it to ranga, he provided the most help, i was the backup :)

OpenStudy (anonymous):

: ( It is saying the answer is wrong.

OpenStudy (anonymous):

But it looks right.

OpenStudy (anonymous):

let me check the math

OpenStudy (anonymous):

does it give you 155√3 + 650/535 instead?

OpenStudy (anonymous):

I don't think so. I think it wants it more simplified or something..

OpenStudy (anonymous):

In the example it gives it is talking about canceling the common factor.

OpenStudy (anonymous):

155√3/535+130/107

OpenStudy (anonymous):

we already did that. but i see the error now. earlier, the answer that was arrived at was 634, not 650. I'm getting 650 now

OpenStudy (anonymous):

It's not wanting to take the 650 either.

OpenStudy (anonymous):

I hate this math program. It is so picky.

OpenStudy (anonymous):

:/ that's strange..I'm not sure how to simplify that further than the two fractions with a common denominator of 535

OpenStudy (anonymous):

Do you want to see if I can send you the example using a screen pic?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Sorry that took so long.

OpenStudy (anonymous):

ohhh i see what they mean. okay

OpenStudy (anonymous):

Are you needing the rest of it?

OpenStudy (anonymous):

no, i think i got it. all the numbers in our solution can be divided by 5. so, 155√3+650/535 would simplify to 31√3+130/107

OpenStudy (anonymous):

5 is the common factor of the solution

OpenStudy (anonymous):

So 31 and all of that all over 107 right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

not a problem! :)

OpenStudy (anonymous):

It's right! Woot!

OpenStudy (anonymous):

:D